THE INFRARED SPECTRUM OF C_{2}H_{4}^{+} TRAPPED IN SOLID NEON

\underline{MARILYN E. JACOX}, WARREN E. THOMPSON, and DANIEL FORNEY, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

When a mixture of C_{2}H_{4} in a large excess of neon is codeposited at 4.3 K with a beam of neon atoms that have been excited in a microwave discharge, the infrared spectrum of the resulting solid includes absorptions of C_{2}H_{4}^{+}. Experiments using isotopically substituted ethylene and density functional calculations support the proposed assignment, which is consistent with recent gas-phase threshold photoelectron spectroscopic observations. Evidence for the stabilization of other products, including the uncharged vinyl radical, will also be presented.